### v1

@@ 7 Experiments
We adopt the pre-trained score function in [30] as our first-order score model. Unfortunately, there is no publicly available second-order score model on the FFHQ dataset. For training the second-order score model, we adopt the NCSN++ architecture [30] that outputs only the diagonal value of the second-order score, yielding the same output dimension as the first-order score model. The focus on the diagonal is due to the fact that only the trace of the second-order score is pertinent to the calculation of the loss function in (21). The first-order score model is adopted to provide the corresponding estimate in (21). The training of the second-order score model follows the same default settings as its first-order counterpart [30]; see Appendix D.1 for more details. For inference, we test our proposed method and several benchmarks on $1000$ images from the validation set of the FFHQ dataset.
(a) $M/N=0.8$, $\beta=1$
(b) $M/N=0.5$, $\beta=0.8$
(c) $M/N=0.2$, $\beta=0.6$
Figure 3: Convergence behaviors (one experiment) of STMP for different compression ratios and noise levels.
Task specification.
@@
We consider a compressive image recovery task with the measurement matrix in the form of $\mathbf{A}=\mathbf{S}\mathbf{W}\bm{\Theta}$, where $\mathbf{S}\in\mathbb{R}^{M\times N}$ is a random row selection matrix consisting of randomly selected rows from a permutation matrix, $\mathbf{W}\in\mathbb{R}^{N\times N}$ is a DCT matrix, and $\bm{\Theta}\in\mathbb{R}^{N\times N}$ is a diagonal matrix with random signs ($1$ or $-1$) in the diagonal. We provide comprehensive results on the image recovery performance under different compression ratios and noise levels.
@@
Ground truth STMP (Ours) DPS DiffPIR PnP-ADMM D-Turbo-CS
@@
Benchmarks.
We compare our proposed method against both the recent state-of-the-art score-based posterior sampling techniques and the traditional plug-and-play methods. The baselines include diffusion posterior sampling (DPS) [6], diffusion models for plug-and-play image restoration (DiffPIR) [36], plug-and-play alternating direction method of multipliers (PnP-ADMM) [4], and denoising-based turbo compressed sensing (D-Turbo-CS) [34]. As suggested in [6, 36], we set the number reverse timesteps (i.e., NFEs) in DPS and DiffPIR as $1000$ and $100$, respectively. We choose the total variation (TV) denoising method as the plug-and-play denoiser in PnP-ADMM, and choose the SURE-LET [3] denoiser for D-Turbo-CS. We provide in Appendix D.2 the hyperparameters used in these benchmarks.
Convergence behaviors and state evolution.
We present the SE prediction only in Fig. 3(a) because the adoption of damping in the other two subfigures makes the SE inaccurate. Overall, the SE equations accurately predict the converged performance of STMP. It is also observed that, before convergence, there is a mismatch between the SE prediction and the actual performance of STMP. We conjecture that this discrepancy stems from the training errors in second-order denoising score matching, leading to a deviation of the estimated posterior variance from Assumption 1.
Qualitative results.
We present some representative results in Fig. 4. Both PnP-ADMM and D-Turbo-CS fail to produce satisfactory results due to the extremely low compression ratio ($M/N=0.1$). When the added noise is small ($\delta_{0}=0.05$), the proposed method, DPS, and DiffPIR are able to generate high-quality reconstructions. The DPS baseline, however, does not always adhere to the ground-truth image when the added noise is large ($\delta_{0}=0.5$). For example, in the last row of Fig. 4, the sample generated by DPS depicts a man wearing transparent glasses, whereas the ground truth shows sunglasses. This highlights the superior restoration fidelity of our method over DPS. More results on different compression ratios and noise levels can be found in Appendix E.
Quantitative results.
In Table 1, we report the average peak signal-to-noise ratio (PSNR), structural similarity index measure (SSIM) [33], Fréchet inception distance (FID) [14], and learned perceptual image patch similarity (LPIPS) [35] to evaluate the faithfulness to the original image and the visual quality. We see that our proposed algorithm outperforms other benchmarks on most of these metrics. More results can be found in Appendix E.
Table 1: Quantitative results on FFHQ validation set. The compression ratio is $M/N=0.1$ and we set $\beta=0.5$.
@@
Performance-complexity tradeoff.
The computational complexity of STMP, DPS, and DiffPIR is dominated by the number of NFEs. In STMP, each iteration involves forward propagation through both the first- and the second-order score networks. Although these two NFEs can be executed in parallel without doubling the time consumption, we still regard each iteration as involving two NFEs. We evaluate the tradeoff between the reconstruction faithfulness/image quality and the computational complexity in Fig. 5. Compared to the other two benchmarks, STMP requires significantly fewer NFEs to produce meaningful results, highlighting the efficiency of our design. Note that DPS requires not only forward propagation during each NFE, but also back-propagating gradients through the score-based network. This makes the sampling steps in DPS even more computationally expensive.
8 Conclusion
@@
In this paper, we introduced score-based turbo message passing (STMP), a novel approach for compressive image recovery using unconditional score-based networks. STMP inherits the fast convergence speed of message passing algorithms and, meanwhile, takes full advantage of the expressive prior through the integration of score-based networks. Theoretically, we demonstrated that the asymptotic performance of STMP can be predicted by a set of scalar equations called state evolution (SE). Experimental results on the FFHQ dataset showcase the state-of-the-art performance of STMP against different benchmarks. This paper lays the groundwork for further enhancing the efficiency of linear inverse problem solving. By making straightforward adjustments to the message calculation in the linear estimation part, STMP can be generalized to address a broader class of inverse problems.
Appendix
Appendix A Proof of Theorem 1
We express the likelihood $p(\tilde{\mathbf{x}}|\mathbf{x})$ in exponential family form as
$p(\tilde{\mathbf{x}}|\mathbf{x})=h(\tilde{\mathbf{x}})\exp\left\{\mathbf{x}^{\sf T}\bm{\varphi}(\tilde{\mathbf{x}})-A(\mathbf{x})\right\},$ (39)
where $h(\tilde{\mathbf{x}})\coloneq\left(2\pi\sigma^{2}\right)^{-\frac{N}{2}}\exp\left\{-\frac{\tilde{\mathbf{x}}^{\sf T}\tilde{\mathbf{x}}}{2\sigma^{2}}\right\}$ is the base measure, $\bm{\varphi}(\tilde{\mathbf{x}})\coloneq\frac{\tilde{\mathbf{x}}}{\sigma^{2}}$ is the sufficient statistic, and $A(\mathbf{x})\coloneq\frac{\mathbf{x}^{\sf T}\mathbf{x}}{2\sigma^{2}}$ is the cumulant function (a.k.a. log-partition function). Eqn. (39) corresponds to the exponential family form of the Gaussian distribution with unknown mean $\mathbf{x}$ and known covariance matrix $\sigma^{2}\mathbf{I}$. Note that since the covariance matrix is assumed known, the sufficient statistic $\bm{\varphi}(\tilde{\mathbf{x}})$ incorporates only the first-order component of $\tilde{\mathbf{x}}$.
By applying the Bayes rule, we can also write the posterior distribution in exponential family form as
$p(\mathbf{x}|\tilde{\mathbf{x}})$ $=\frac{p(\mathbf{x})h(\tilde{\mathbf{x}})\exp\left\{\mathbf{x}^{\sf T}\bm{\varphi}(\tilde{\mathbf{x}})-A(\mathbf{x})\right\}}{p(\tilde{\mathbf{x}})}$ (40)
$=\bar{h}(\mathbf{x})\exp\left\{\tilde{\mathbf{x}}^{\sf T}\bar{\bm{\varphi}}(\mathbf{x})-\bar{A}(\tilde{\mathbf{x}})\right\},$ (41)
where $\bar{h}(\mathbf{x})\coloneq p(\mathbf{x})\exp\left\{-A(\mathbf{x})\right\}$, $\bar{\bm{\varphi}}(\mathbf{x})\coloneq\frac{\mathbf{x}}{\sigma^{2}}$, and $\bar{A}(\tilde{\mathbf{x}})\coloneq\log\frac{p(\tilde{\mathbf{x}})}{h(\tilde{\mathbf{x}})}$.
For exponential families, the mean and covariance matrix of the sufficient statistics can be calculated by taking the first- and second-order derivatives of the cumulant function, respectively [2]. That is,
$\mathbb{E}\left[\bar{\bm{\varphi}}(\mathbf{x})|\tilde{\mathbf{x}}\right]$ $=\nabla_{\tilde{\mathbf{x}}}\bar{A}(\tilde{\mathbf{x}})=\nabla_{\tilde{\mathbf{x}}}\log p(\tilde{\mathbf{x}})-\nabla_{\tilde{\mathbf{x}}}\log h(\tilde{\mathbf{x}}),$ (42)
$\mathrm{Cov}\left[\bar{\bm{\varphi}}(\mathbf{x})|\tilde{\mathbf{x}}\right]$ $=\nabla_{\tilde{\mathbf{x}}}^{2}\bar{A}(\tilde{\mathbf{x}})=\nabla_{\tilde{\mathbf{x}}}^{2}\log p(\tilde{\mathbf{x}})-\nabla_{\tilde{\mathbf{x}}}^{2}\log h(\tilde{\mathbf{x}}).$ (43)
Similar relationships hold for higher-order moments of the sufficient statistics. Substituting $\bar{\bm{\varphi}}(\mathbf{x})=\frac{\mathbf{x}}{\sigma^{2}}$ and $h(\tilde{\mathbf{x}})=\left(2\pi\sigma^{2}\right)^{-\frac{N}{2}}\exp\left\{-\frac{\tilde{\mathbf{x}}^{\sf T}\tilde{\mathbf{x}}}{2\sigma^{2}}\right\}$ into (42) and (43) yields the desired result in Theorem 1.
Appendix B Derivation of the Second-Order Denoising Score Matching
We provide in Lemma 1 an expectation formulation of the first-order score function.
Lemma 1.
For $(\mathbf{x},\tilde{\mathbf{x}})\sim p(\mathbf{x},\tilde{\mathbf{x}})$, we have
$\nabla_{\tilde{\mathbf{x}}}\log p(\tilde{\mathbf{x}})=\mathbb{E}_{p(\mathbf{x}|\tilde{\mathbf{x}})}\left[\nabla_{\tilde{\mathbf{x}}}\log p(\tilde{\mathbf{x}}|\mathbf{x})\right].$ (44)
Proof.
$\nabla_{\tilde{\mathbf{x}}}\log p(\tilde{\mathbf{x}})$ $=\frac{\nabla_{\tilde{\mathbf{x}}}p(\tilde{\mathbf{x}})}{p(\tilde{\mathbf{x}})}$ (45)
$=\frac{\nabla_{\tilde{\mathbf{x}}}\int p(\mathbf{x})p(\tilde{\mathbf{x}}|\mathbf{x})\mathrm{d}\mathbf{x}}{p(\tilde{\mathbf{x}})}$ (46)
$=\int\frac{p(\mathbf{x})p(\tilde{\mathbf{x}}|\mathbf{x})}{p(\tilde{\mathbf{x}})}\frac{\nabla_{\tilde{\mathbf{x}}}p(\tilde{\mathbf{x}}|\mathbf{x})}{p(\tilde{\mathbf{x}}|\mathbf{x})}\mathrm{d}\mathbf{x}$ (47)
$=\int p(\mathbf{x}|\tilde{\mathbf{x}})\nabla_{\tilde{\mathbf{x}}}\log p(\tilde{\mathbf{x}}|\mathbf{x})\mathrm{d}\mathbf{x}$ (48)
$=\mathbb{E}_{p(\mathbf{x}|\tilde{\mathbf{x}})}\left[\nabla_{\tilde{\mathbf{x}}}\log p(\tilde{\mathbf{x}}|\mathbf{x})\right].$ (49)
We provide in Lemma 2 an expectation formulation of the second-order score function.
Lemma 2.
For $(\mathbf{x},\tilde{\mathbf{x}})\sim p(\mathbf{x},\tilde{\mathbf{x}})$, we have
$\nabla_{\tilde{\mathbf{x}}}^{2}\log p(\tilde{\mathbf{x}})=\mathbb{E}_{p(\mathbf{x}|\tilde{\mathbf{x}})}\left[\nabla_{\tilde{\mathbf{x}}}^{2}\log p(\tilde{\mathbf{x}}|\mathbf{x})+\mathbf{b}(\mathbf{x},\tilde{\mathbf{x}})\mathbf{b}(\mathbf{x},\tilde{\mathbf{x}})^{\sf T}\right],$ (50)
where $\mathbf{b}(\mathbf{x},\tilde{\mathbf{x}})\coloneq\nabla_{\tilde{\mathbf{x}}}\log p(\tilde{\mathbf{x}})-\nabla_{\tilde{\mathbf{x}}}\log p(\tilde{\mathbf{x}}|\mathbf{x})$.
Proof.
Firstly, the gradient of $p(\mathbf{x}|\tilde{\mathbf{x}})$ w.r.t. $\tilde{\mathbf{x}}$ can be calculated as
$\nabla_{\tilde{\mathbf{x}}}p(\mathbf{x}|\tilde{\mathbf{x}})$ $=\nabla_{\tilde{\mathbf{x}}}\frac{p(\mathbf{x})p(\tilde{\mathbf{x}}|\mathbf{x})}{p(\tilde{\mathbf{x}})}$ (51)
$=p(\mathbf{x})\frac{p(\tilde{\mathbf{x}})\nabla_{\tilde{\mathbf{x}}}p(\tilde{\mathbf{x}}|\mathbf{x})-p(\tilde{\mathbf{x}}|\mathbf{x})\nabla_{\tilde{\mathbf{x}}}p(\tilde{\mathbf{x}})}{p(\tilde{\mathbf{x}})^{2}}$ (52)
$=\frac{p(\mathbf{x})p(\tilde{\mathbf{x}}|\mathbf{x})}{p(\tilde{\mathbf{x}})}\left(\frac{\nabla_{\tilde{\mathbf{x}}}p(\tilde{\mathbf{x}}|\mathbf{x})}{p(\tilde{\mathbf{x}}|\mathbf{x})}-\frac{\nabla_{\tilde{\mathbf{x}}}p(\tilde{\mathbf{x}})}{p(\tilde{\mathbf{x}})}\right)$ (53)
$=p(\mathbf{x}|\tilde{\mathbf{x}})\left(\nabla_{\tilde{\mathbf{x}}}\log p(\tilde{\mathbf{x}}|\mathbf{x})-\nabla_{\tilde{\mathbf{x}}}\log p(\tilde{\mathbf{x}})\right).$ (54)
Then, we apply Lemma 1 to the second-order score function:
$\!\!\!\!\nabla_{\tilde{\mathbf{x}}}^{2}\log p(\tilde{\mathbf{x}})$ $=\nabla_{\tilde{\mathbf{x}}}\mathbb{E}_{p(\mathbf{x}|\tilde{\mathbf{x}})}\left[\nabla_{\tilde{\mathbf{x}}}\log p(\tilde{\mathbf{x}}|\mathbf{x})\right]$ (55)
$=\int\Big{(}p(\mathbf{x}|\tilde{\mathbf{x}})\nabla_{\tilde{\mathbf{x}}}^{2}\log p(\tilde{\mathbf{x}}|\mathbf{x})+\nabla_{\tilde{\mathbf{x}}}p(\mathbf{x}|\tilde{\mathbf{x}})\nabla_{\tilde{\mathbf{x}}}\log p(\tilde{\mathbf{x}}|\mathbf{x})^{\sf T}\Big{)}\mathrm{d}\mathbf{x}$ (56)
$=\mathbb{E}_{p(\mathbf{x}|\tilde{\mathbf{x}})}\Big{[}\nabla_{\tilde{\mathbf{x}}}^{2}\log p(\tilde{\mathbf{x}}|\mathbf{x})+\big{(}\nabla_{\tilde{\mathbf{x}}}\log p(\tilde{\mathbf{x}}|\mathbf{x})-\nabla_{\tilde{\mathbf{x}}}\log p(\tilde{\mathbf{x}})\big{)}\nabla_{\tilde{\mathbf{x}}}\log p(\tilde{\mathbf{x}}|\mathbf{x})^{\sf T}\Big{]},$ (57)
$=\mathbb{E}_{p(\mathbf{x}|\tilde{\mathbf{x}})}\Big{[}\nabla_{\tilde{\mathbf{x}}}^{2}\log p(\tilde{\mathbf{x}}|\mathbf{x})-\mathbf{b}(\mathbf{x},\tilde{\mathbf{x}})\nabla_{\tilde{\mathbf{x}}}\log p(\tilde{\mathbf{x}}|\mathbf{x})^{\sf T}\Big{]},$ (58)
where (57) applies the result in (54). From Lemma 1, we also have
$\mathbb{E}_{p(\mathbf{x}|\tilde{\mathbf{x}})}\left[\mathbf{b}(\mathbf{x},\tilde{\mathbf{x}})\right]$ $=\mathbb{E}_{p(\mathbf{x}|\tilde{\mathbf{x}})}\left[\nabla_{\tilde{\mathbf{x}}}\log p(\tilde{\mathbf{x}})-\nabla_{\tilde{\mathbf{x}}}\log p(\tilde{\mathbf{x}}|\mathbf{x})\right]$ (59)
$=\nabla_{\tilde{\mathbf{x}}}\log p(\tilde{\mathbf{x}})-\mathbb{E}_{p(\mathbf{x}|\tilde{\mathbf{x}})}\left[\nabla_{\tilde{\mathbf{x}}}\log p(\tilde{\mathbf{x}}|\mathbf{x})\right]$ (60)
$=\mathbf{0},$ (61)
and therefore
$\mathbb{E}_{p(\mathbf{x}|\tilde{\mathbf{x}})}\left[\mathbf{b}(\mathbf{x},\tilde{\mathbf{x}})\nabla_{\tilde{\mathbf{x}}}\log p(\tilde{\mathbf{x}})^{\sf T}\right]=\mathbb{E}_{p(\mathbf{x}|\tilde{\mathbf{x}})}\left[\mathbf{b}(\mathbf{x},\tilde{\mathbf{x}})\right]\nabla_{\tilde{\mathbf{x}}}\log p(\tilde{\mathbf{x}})^{\sf T}=\mathbf{0}.$ (62)
Thus, we can further transform (58) by adding $\mathbf{0}$ as
$\nabla_{\tilde{\mathbf{x}}}^{2}\log p(\tilde{\mathbf{x}})$ $=\mathbb{E}_{p(\mathbf{x}|\tilde{\mathbf{x}})}\left[\nabla_{\tilde{\mathbf{x}}}^{2}\log p(\tilde{\mathbf{x}}|\mathbf{x})-\mathbf{b}(\mathbf{x},\tilde{\mathbf{x}})\nabla_{\tilde{\mathbf{x}}}\log p(\tilde{\mathbf{x}}|\mathbf{x})^{\sf T}\right]+\mathbb{E}_{p(\mathbf{x}|\tilde{\mathbf{x}})}\left[\mathbf{b}(\mathbf{x},\tilde{\mathbf{x}})\nabla_{\tilde{\mathbf{x}}}\log p(\tilde{\mathbf{x}})^{\sf T}\right]$ (63)
$=\mathbb{E}_{p(\mathbf{x}|\tilde{\mathbf{x}})}\left[\nabla_{\tilde{\mathbf{x}}}^{2}\log p(\tilde{\mathbf{x}}|\mathbf{x})+\mathbf{b}(\mathbf{x},\tilde{\mathbf{x}})\mathbf{b}(\mathbf{x},\tilde{\mathbf{x}})^{\sf T}\right],$ (64)
which completes the proof of Lemma 2. ∎
By noting that $\nabla_{\tilde{\mathbf{x}}}\log p(\tilde{\mathbf{x}}|\mathbf{x})=-\frac{\tilde{\mathbf{x}}-\mathbf{x}}{\sigma^{2}}$ and $\nabla_{\tilde{\mathbf{x}}}^{2}\log p(\tilde{\mathbf{x}}|\mathbf{x})=-\frac{\mathbf{I}}{\sigma^{2}}$, we rewrite (18) as
$\min_{\bm{\phi}}~{}\mathbb{E}_{p(\tilde{\mathbf{x}})}\mathbb{E}_{p(\mathbf{x}|\tilde{\mathbf{x}})}\left[\left\|\mathbf{S}_{\bm{\phi}}(\tilde{\mathbf{x}},\sigma)-\mathbf{b}(\mathbf{x},\tilde{\mathbf{x}},\sigma)\mathbf{b}(\mathbf{x},\tilde{\mathbf{x}},\sigma)^{\sf T}-\nabla_{\tilde{\mathbf{x}}}^{2}\log p(\tilde{\mathbf{x}}|\mathbf{x})\right\|^{2}_{F}\right].$ (65)
Note that the dependence on $\sigma$ in $\mathbf{b}(\mathbf{x},\tilde{\mathbf{x}},\sigma)$ is sometimes omitted to simplify the notation. In (65), minimizing the inner expectation is an MMSE problem, whose optimal solution satisfies
$\mathbf{S}_{\bm{\phi}^{\star}}(\tilde{\mathbf{x}},\sigma)=\mathbb{E}_{p(\mathbf{x}|\tilde{\mathbf{x}})}\left[\mathbf{b}(\mathbf{x},\tilde{\mathbf{x}},\sigma)\mathbf{b}(\mathbf{x},\tilde{\mathbf{x}},\sigma)^{\sf T}+\nabla_{\tilde{\mathbf{x}}}^{2}\log p(\tilde{\mathbf{x}}|\mathbf{x})\right].$ (66)
Applying Lemma 2 to (66) yields
$\mathbf{S}_{\bm{\phi}^{\star}}(\tilde{\mathbf{x}},\sigma)=\nabla_{\tilde{\mathbf{x}}}^{2}\log p(\tilde{\mathbf{x}}),$ (67)
which completes the proof.
Appendix C Supplementary Information for State Evolution Analysis
C.1 Additional Assumptions
Assumption 1 (Right rotationally invariant).
The distribution of $\mathbf{A}$ is equal to that of $\mathbf{A}\mathbf{V}$ for any orthogonal matrix $\mathbf{V}$, independent of $\mathbf{A}$.
Assumption 2 (Uniformly Lipschitz continuous).
The denoiser $\mathbf{g}_{\sf B}(\cdot)$ is uniformly Lipschitz continuous with parameters $\alpha_{1}$, $\alpha_{2}$, $\alpha_{3}$, i.e.,
$\left\|\mathbf{g}_{\sf B}(\mathbf{r}_{2},v_{2})-\mathbf{g}_{\sf B}(\mathbf{r}_{1},v_{1})\right\|\leq(\alpha_{1}+\alpha_{2}|v_{2}-v_{1}|)\left\|\mathbf{r}_{2}-\mathbf{r}_{1}\right\|+\alpha_{3}\sqrt{N}|v_{2}-v_{1}|,$ (68)
for any $\mathbf{r}_{1}$, $\mathbf{r}_{2}$, $v_{1}$, $v_{2}$, and $N$.
Assumption 3 (Convergent under Gaussian noise).
The vector $\mathbf{x}\in\mathbb{R}^{N}$ and the denoiser $\mathbf{g}_{\sf B}(\cdot)$ are convergent under Gaussian noise. Specifically, let $\mathbf{w}_{1},\mathbf{w}_{2}\in\mathbb{R}^{N}$ be two vectors where $(w_{1n},w_{2n})$ are i.i.d. with $(w_{1n},w_{2n})=\mathcal{N}(\mathbf{0},\mathbf{S})$ for some positive definite covariance $\mathbf{S}\in\mathbb{R}^{2\times 2}$. The following limits exist almost surely:
$\lim_{N\rightarrow\infty}\frac{1}{N}\mathbf{g}_{\sf B}(\mathbf{x}+\mathbf{w}_{1},v_{1})^{\sf T}\mathbf{g}_{\sf B}(\mathbf{x}+\mathbf{w}_{2},v_{2}),~{}~{}~{}~{}\lim_{N\rightarrow\infty}\frac{1}{N}\mathbf{g}_{\sf B}(\mathbf{x}+\mathbf{z}_{1},v_{1})^{\sf T}\mathbf{x},$ (69)
$\lim_{N\rightarrow\infty}\frac{1}{N}\mathbf{x}^{\sf T}\mathbf{w}_{1},~{}~{}~{}~{}\lim_{N\rightarrow\infty}\frac{1}{N}\left\|\mathbf{x}\right\|^{2}_{2},$ (70)
$\lim_{N\rightarrow\infty}\frac{1}{N}\mathbf{g}(\mathbf{x}+\mathbf{w}_{1},v_{1})^{\sf T}\mathbf{w}_{2},$ (71)
for all $v_{1}$, $v_{2}$, and covariance matrices $\mathbf{S}$. Moreover, the values of the limits are continuous in $\mathbf{S}$, $v_{1}$, and $v_{2}$.
C.2 Further Clarifications
Although Turbo-CS [21] and VAMP [13, 25] share the same idea, they use different definitions for their variables. This paper adopts the notations used in Turbo-CS. To highlight the equivalence, we summarize in Table 2 the relationship between the SE variables used in this paper and these used in [13]. With this one-to-one correspondence, on can verify the correctness of Theorem 2 by referring to [13], which provides a rigorous proof of SE using VAMP’s notation.
Table 2: One-to-one correspondence between the SE variables in this paper and these in [13].
Variable in [13] Variable in this paper Variable in [13] Variable in this paper
$\bar{\alpha}_{1k}$ $v_{{\sf B},k}^{\textit{post}}/v_{{\sf B},k}^{\textit{pri}}$ $\bar{\alpha}_{2k}$ $v_{{\sf A},k}^{\textit{post}}/v_{{\sf A},k}^{\textit{pri}}$
$\bar{\gamma}_{1k}$ $1/v_{{\sf B},k}^{\textit{pri}}$ $\bar{\gamma}_{2k}$ $1/v_{{\sf A},k}^{\textit{pri}}$
$\bar{\eta}_{1k}$ $1/v_{{\sf B},k}^{\textit{post}}$ $\bar{\eta}_{2k}$ $1/v_{{\sf A},k}^{\textit{post}}$
$\tau_{1k}$ $\tau_{{\sf B},k}$ $\tau_{2k}$ $\tau_{{\sf A},k}$
Appendix D Training and Inference Details
D.1 Training Configurations
We use the pretrained checkpoint in [30] for the first-order score network. For the second-order score network, we use the same NCSN++ architecture in [30]. It outputs only the diagonal value of the second-order score, yielding the same output dimension as the first-order score network. As suggested by [19], the weighting factor in the training objective (21) is chosen as $\lambda_{2}(\sigma_{i})=1/\sigma_{i}^{4}$. As a side note, for the first-order score matching, the work in [30] uses the weighting factor $\lambda_{1}(\sigma_{i})=1/\sigma_{i}^{2}$, which is consistent with the suggestion in [19]. For the second-order score matching, we adopt the SDE training framework in [30] that applies to continuous noise scales in the range of $[0.01,348]$. We use a batch size of $64$. The other training parameters are set identically to its first-order counterpart [30]. We train the second-order score network for $1$M iterations using $4$ NVIDIA A800 GPUs.
Code and checkpoints will be open-sourced upon the acceptance of this paper.
D.2 Inference Configurations
For inference, the proposed method, DPS, and DiffPIR all run on a single NVIDIA A800 GPU. We tune the hyperparameters of the benchmarks as follows:
  • DPS [6]: We use $1000$ diffusion steps by default. We choose $\rho=1$ for the best performance. For obtaining the tradeoff curves in Fig. 5, we also test $50$, $100$, $200$, $400$ reverse diffusion steps.
  • DiffPIR [36]: We use $100$ diffusion steps by default. We choose $\lambda=5$ and $\zeta=1$ for the best performance. For obtaining the tradeoff curves in Fig. 5, we also test $10$, $20$, $40$, $60$ reverse diffusion steps.
  • PnP-ADMM [4]: We choose the total variation (TV) denoiser for PnP-ADMM. We choose $\lambda=0.01$, $\rho=1$, and $\gamma=1$ for the best performance. The algorithm is executed for $50$ iterations.
  • D-Turbo-CS [34]: We choose the SURE-LET [3] denoiser for D-Turbo-CS. There is no hyperparameters for D-Turbo-CS. The algorithm is executed for $50$ iterations.
Appendix E More Results
We provide additional quantitative results in Tables 3 and 4. More recovered images under different compression ratios and noise levels can be found in Figs. 6, 7, and 8.
Table 3: Quantitative results on FFHQ validation set. The compression ratio is $M/N=0.4$ and we set $\beta=0.8$.
Method $\delta_{0}=0.5$ $\delta_{0}=0.05$
[…]
STMP (Ours) $\mathbf{24.73}$ $\mathbf{0.7320}$ $61.22$ $\mathbf{0.0359}$ $\mathbf{33.49}$ $\mathbf{0.9249}$ $31.35$ $\mathbf{0.0040}$
DPS $16.75$ $0.5466$ $\mathbf{44.89}$ $0.0830$ $26.66$ $\underline{0.8273}$ $\underline{26.46}$ $0.0116$
DiffPIR $\underline{23.18}$ $\underline{0.6461}$ $\underline{59.82}$ $\underline{0.0464}$ $28.23$ $0.8100$ $\mathbf{24.22}$ $\underline{0.008}1$
PnP-ADMM $8.75$ $0.0381$ $352.42$ $0.7221$ $\underline{28.89}$ $0.7530$ $56.15$ $0.0118$
D-Turbo-CS $8.62$ $0.0088$ $372.86$ $0.8048$ $24.77$ $0.5723$ $117.74$ $0.0340$
Table 4: Quantitative results on FFHQ validation set. The compression ratio is $M/N=0.7$ and we set $\beta=1$.
Method $\delta_{0}=0.5$ $\delta_{0}=0.05$
[…]
STMP (Ours) $\mathbf{25.82}$ $\mathbf{0.7649}$ $\underline{57.76}$ $\mathbf{0.0280}$ $\mathbf{34.98}$ $\mathbf{0.9420}$ $\mathbf{23.42}$ $\mathbf{0.0027}$
DPS $17.98$ $0.5888$ $\mathbf{39.98}$ $0.0653$ $28.33$ $\underline{0.8596}$ $\underline{26.63}$ $\underline{0.0081}$
DiffPIR $\underline{24.19}$ $\underline{0.7015}$ $86.51$ $\underline{0.0407}$ $28.05$ $0.7728$ $27.00$ $0.0089$
PnP-ADMM $8.68$ $0.0400$ $341.78$ $0.7623$ $\underline{29.15}$ $0.7314$ $43.26$ $0.0101$
D-Turbo-CS $13.92$ $0.0234$ $352.49$ $0.5016$ $28.04$ $0.7387$ $74.22$ $0.0167$
Ground truth STMP (Ours) DPS DiffPIR PnP-ADMM D-Turbo-CS
Figure 6: More results on FFHQ validation set. The compression ratio is $M/N=0.1$ and we set $\beta=0.5$.
Ground truth STMP (Ours) DPS DiffPIR PnP-ADMM D-Turbo-CS
Figure 7: More results on FFHQ validation set. The compression ratio is $M/N=0.4$ and we set $\beta=0.8$.
Ground truth STMP (Ours) DPS DiffPIR PnP-ADMM D-Turbo-CS
Figure 8: More results on FFHQ validation set. The compression ratio is $M/N=0.7$ and we set $\beta=1$.

### v2

@@ VI Experiments
In this section, we evaluate the performance of the proposed method based on the FFHQ $256\times 256$ dataset [14]. We adopt the pre-trained score function in [8] as our first-order score model. Unfortunately, there is no publicly available second-order score model on the FFHQ dataset. For training the second-order score model, we adopt the NCSN++ architecture [8] that outputs only the diagonal value of the second-order score, yielding the same output dimension as the first-order score model. The training of the second-order score model follows the same default settings as its first-order counterpart [8]. For inference, we test our proposed method and several benchmarks on $1000$ images from the validation set of the FFHQ dataset.
@@
We consider a compressive image recovery task with the measurement matrix in the form of $\mathbf{A}=\mathbf{S}\mathbf{W}\boldsymbol{\Theta}$, where $\mathbf{S}\in\mathbb{R}^{M\times N}$ is a random row selection matrix consisting of randomly selected rows from a permutation matrix, $\mathbf{W}\in\mathbb{R}^{N\times N}$ is a DCT matrix, and $\boldsymbol{\Theta}\in\mathbb{R}^{N\times N}$ is a diagonal matrix with random signs ($1$ or 1-1) in the diagonal. We provide comprehensive results on the image recovery performance under different compression ratios and noise levels.
We compare our proposed method against both the recent state-of-the-art score-based posterior sampling techniques and the traditional plug-and-play methods. The baselines include diffusion posterior sampling (DPS) [10], diffusion models for plug-and-play image restoration (DiffPIR) [11], plug-and-play alternating direction method of multipliers (PnP-ADMM) [20], and D-Turbo-CS [3]. As suggested in [10, 11], we set the number reverse timesteps (i.e., NFEs) in DPS and DiffPIR as $1000$ and $100$, respectively. We choose the total variation (TV) denoising method as the plug-and-play denoiser in PnP-ADMM, and choose the SURE-LET [4] denoiser for D-Turbo-CS.
(a) $M/N=0.5$, $\beta=0.8$
(b) $M/N=0.2$, $\beta=0.6$
Figure 2: Convergence behaviors of STMP for different compression ratios and noise levels.
@@
Ground truth STMP DPS DiffPIR PnP-ADMM D-Turbo-CS
@@
We present some representative results in Fig. 3. Both PnP-ADMM and D-Turbo-CS fail to produce satisfactory results due to the extremely low compression ratio ($M/N=0.1$). When the added noise is small ($\delta_{0}=0.05$), the proposed method, DPS, and DiffPIR are able to generate high-quality reconstructions. The DPS baseline, however, does not always adhere to the ground-truth image when the added noise is large ($\delta_{0}=0.5$). For example, in the last row of Fig. 3, the sample generated by DPS depicts a man wearing transparent glasses, whereas the ground truth shows sunglasses. This highlights the superior restoration fidelity of our method over DPS. In Table I, we report the average peak signal-to-noise ratio (PSNR), structural similarity index measure (SSIM), Fréchet inception distance (FID), and learned perceptual image patch similarity (LPIPS) to evaluate the faithfulness to the original image and the visual quality. We see that our proposed algorithm outperforms other benchmarks on most of these metrics.
TABLE I: Quantitative results on FFHQ validation set. The compression ratio is $M/N=0.1$ and we set $\beta=0.5$.
@@
The computational complexity of STMP, DPS, and DiffPIR is dominated by the number of NFEs. In STMP, each iteration involves forward propagation through both the first- and the second-order score networks. Although these two NFEs can be executed in parallel without doubling the time consumption, we still consider each iteration as two NFEs. We evaluate the tradeoff between the reconstruction faithfulness/image quality and the computational complexity in Fig. 4. Compared to the other two benchmarks, STMP requires significantly fewer NFEs to produce meaningful results, highlighting the efficiency of our design. Note that DPS requires not only forward propagation during each NFE, but also back-propagating gradients through the score network. This makes the sampling steps in DPS even more computationally expensive.
@@
VII Conclusion
@@
In this paper, we introduced score-based turbo message passing (STMP), a novel approach for compressive image recovery relying on score-based generative modeling. STMP inherits the fast convergence speed of message passing algorithms and, meanwhile, takes full advantage of the expressive prior through the integration of score networks. Experimental results on the FFHQ dataset showcase the state-of-the-art performance of STMP against different benchmarks.